# Navigating the Lobbying Landscape: Insights from Opinion Dynamics Models

### v2

@@ Abstract
While lobbying has been demonstrated to have an important effect on public opinion and policy making, existing models of opinion formation do not specifically include its effect. In this work we introduce a new model of opinion dynamics where lobbyists can implement complex strategies and are characterised by a finite budget. Individuals update their opinions through a learning process resembling Bayesian learning, but influenced by cognitive biases such as under-reaction and confirmation bias. We study the model numerically and demonstrate rich dynamics both with and without lobbyists. In the presence of lobbying, we observe two regimes: one in which lobbyists can have full influence on the agent network, and another where the peer-effect generates polarisation. When symmetric lobbyists are present, the lobbyist influence regime is characterised by long opinion oscillations, while in the transition area between the two regimes we observe convergence to the optimistic model when the lobbying influence is long enough. These rich dynamics pave the way for studying real lobbying strategies to validate the model in practice.
@@ Introduction
This paper contributes to this growing interdisciplinary literature by introducing a novel model of lobbying influence over opinion dynamics in a boundedly rational social network. In our framework, a population of agents interacts repeatedly to update their subjective beliefs about the likelihood of an uncertain future event, for example, a damaging climate event. Each agent weighs two competing probabilistic models—one “optimistic” and one “pessimistic”—and updates their beliefs through Bayesian-like learning, modulated by under-reaction[24, 25, 26] and confirmation bias (a form of directional motivated reasoning) [27]. Agents are embedded in a directed network and interact through sequential signal exchanges that propagate beliefs over time. We then endogenize lobbying by introducing external agents — lobbyists — who strategically send costly signals to influence the population’s belief distribution by the end of a finite time horizon. Lobbyists differ in the model they support and are subject to a resource constraint (budget). Their objective is to minimize the distance between the final average belief and their preferred model. Lobbyists select randomized strategies from a constrained set of signaling paths, and their interventions shape both the speed and direction of belief evolution in the network.
@@
Our numerical results reveal rich dynamics. In the absence of lobbying, the network tends to converge toward consensus or, under high confirmation bias, become polarized. A single lobbyist with sufficient resources can dominate belief formation and steer the network toward its preferred model, especially when agents are relatively open to new information. However, when two lobbyists with opposing goals act simultaneously, the system often fails to converge within the active time horizon, resulting in persistent instability and oscillation — especially in regions of the parameter space where both lobbyists are highly effective. We also uncover non-trivial path dependence and sensitivity to initial conditions, pointing to the complex interplay between cognitive constraints, network structure, and strategic influence.
@@
The remainder of the paper is structured as follows. The next section introduces the baseline opinion dynamics model, describing the agent interaction protocol and belief updating mechanism, and then incorporates lobbying agents and formalizes their strategic behavior, cost constraints, and impact on individual beliefs. The Results section presents the simulation results across three scenarios: no lobbying, single-lobbyist, and dual-lobbyist configurations. It explores the model’s behavior across key parameters (underreaction and confirmation bias), and performs sensitivity analysis on budget, network size and time horizon. We then conclude by summarizing the findings and suggesting avenues for future research.
@@
Consider $N$ individuals who interact for $\tau$ rounds about the occurrence of an uncertain damaging event after the interaction rounds. Individuals do not know the true probabilities that determine the occurrence of the event, but can rely on two probabilistic models. Those are the probability distributions $(\pi_{o},1-\pi_{o})$ and $(\pi_{p},1-\pi_{p})$, where $\pi_{k}$, with $k\in\{o,p\}$, indicates the probability that the damaging event occurs. The model $(\pi_{o},1-\pi_{o})$ is optimistic while the model $(\pi_{p},1-\pi_{p})$ is pessimistic: $10<\pi_{o}<\pi_{p}<$ 1 with $\pi_{o}$ close to zero and $\pi_{p}$ close to one.
@@
where $\lambda_{i,t}\in[0,1]$ modulates the effects of cognitive biases of agent $i$ at round $t$ and $w_{i,0}\in(0,1)$ is the initial prior. If $\lambda_{i,t}=0$ then $i$ is a Bayesian learner and updates probabilities according to Bayes rule[28]. If $\lambda_{i,t}=1$ then $i$ keeps the initial combination fixed no matter the sequence of signals it receives. Intermediate values of $\lambda_{i,t}$ indicate that, as $i$ receives a signal, it updates the weights in the right direction but in lower magnitude than what Bayesian learning would prescribe. Thus, individuals under-react to information [24, 25, 26] and we enrich such a behavioural interpretation imposing that $\lambda_{i,t}$ depends upon the type of signal received and the prior. In this way, it captures a form of directional motivated reasoning that, in turn, gives rise to confirmation bias (the tendency of individuals to discard information that contradicts their priors), a common feature in the formation of beliefs [27]. In particular, we assume
@@
with $\phi_{i},\lambda_{i}\in[0,1]$. That is, the overall degree of under-reaction is the convex combination of two components. On the one hand, the parameter $\lambda_{i}$ captures the baseline level of under-reaction of agent $i$. On the other hand, the function $|1-s_{t}-w_{i,t-1}|$ introduces directional motivated reasoning or, equivalently for our analysis and interpretation, confirmation bias. The parameter $\phi_{i}$ regulates the strength of the directional motivated reasoning. Suppose, for instance, that directional motivated reasoning is strong ($\phi_{i}\simeq 1$), if an individual has a strong prior in the optimistic model ($w_{i,t-1}\simeq 1$) and receives a signal that favours it with respect to the pessimistic model ($s_{t}=0$) then its overall degree of under-reaction is low: information has a sensible impact on its beliefs. If, instead, the information contradicts its priors ($s_{t}=1$), then the degree of under-reaction is high: information has a negligible effect on its beliefs. A symmetric argument holds if the agent’s priors favour the pessimistic model. If, instead, directional motivated reasoning is weak ($\phi_{i}\simeq 0$), then the individual under-reacts with respect to information because of the effect of $\lambda_{i}$, but the nature of the signal and its relation to the prior have a negligible effect on probability updating.
Assuming that the signal at round $t$ is drawn from the distribution ${\bf p}_{\iota_{t},t}$ (i.e., $\text{Prob}\{s_{t}=1\}=p_{\iota_{t},t}$), the timeline of events in each round $t$ is as follows: $i)$ $\iota_{t}$ is independently and uniformly drawn from the population; $ii)$ the set of receivers $J(\iota_{t})$ is created; $iii)$ the signal $s_{t}$ is drawn from the distribution ${\bf p}_{\iota_{t},t}$; $iv)$ individual weights are updated according to Eq. (2).
@@
Lobbyists can be considered external agents with respect to the social network that, in each turn, can pay a cost to send signals to individuals. Their objective is to influence the time $\tau$ distribution of beliefs so that it becomes concentrated on one of the two models.
@@
Assume that there are $L$ active lobbyists and define $S_{i,t}^{\ell}\in\{0,1\}$ as a variable indicating whether lobbyist $\ell\in\{1,2,\ldots,L\}$ sends a signal to individual $i$ at the beginning of round $t$ ($S^{\ell}_{i,t}=1$) or not ($S^{\ell}_{i,t}=0$). It follows that a pure strategy for lobbyist $\ell$ can be indicated as a matrix ${S}^{\ell}\in\mathcal{S}=\{0,1\}^{N\times\tau}$. Each lobbyist supports one model between the optimist and the pessimist one and the signals it sends are in favour of the model it supports. For instance, a lobbyist supporting the optimistic model will signal 0, that is, the event is not occurring. Define the function that assigns to each lobbyist the model it supports as $m:\{1,2,\ldots,L\}\to\{o,p\},\ell\mapsto m(\ell)$. We assume that signals are costly. In particular, the cost of a signal is fixed to a given level $c>0$ and each lobbyist $\ell$ has a budget $B^{\ell}$. Without loss of generality, we shall set $c=1$ in what follows. That is, $B^{\ell}$ shall be understood as the maximum number of signals lobbyist $\ell$ can send. Notice that this assumption constrains the set of strategies that each lobbyist can actually play, that is, a feasible strategy is such that $\sum_{t=1}^{\tau}\sum_{i=1}^{N}\,S^{\ell}_{i,t}\leq B^{\ell}$. Hence, we define the set of feasible strategies for lobbyist $\ell$ as $\mathcal{S}_{\ell}=\{S^{\ell}\in\mathcal{S}\,|\,\sum_{t=1}^{\tau}\sum_{i=1}^{N}\,S^{\ell}_{i,t}\leq B^{\ell}\}$. We allow lobbyists to randomize their choices. Call $\Delta(\mathcal{S}_{\ell})$ the mixed extension of $\mathcal{S}_{\ell}$ and $\sigma^{\ell}({S})$ the probability that lobbyist $\ell$ attaches to a feasible strategy ${S}$, such that $\sigma^{\ell}\in\Delta(\mathcal{S}_{\ell})$ is a feasible mixed strategy for lobbyist $\ell$.
Each individual reacts to the signal received in the same way as they do for peer signals. Therefore, now, individual probabilities also evolve as a consequence of the strategies of the lobbyists. In particular, call $S^{-\ell}=(S^{1},\ldots,S^{\ell-1},S^{\ell+1},\ldots,S^{L})\in\mathcal{S}_{1}\times\ldots\times\mathcal{S}_{\ell-1}\times\mathcal{S}_{\ell+1}\times\ldots\times\mathcal{S}_{L}$ a profile of feasible pure strategies for the lobbyists different from $\ell$, such that $p_{i,\tau}(S^{\ell},S^{-\ell})$ indicates the final probability of individual $i$ as a function of the strategies played by all lobbyists. Define the vector of lobbyists’ signals received by agent $i$ at time $t$ as $\mathcal{L}^{i,t}\in\{0,1,2\}^{L}$, where $\mathcal{L}^{i,t}_{\ell}=0$ if $S^{\ell}_{i,t}=0$, $\mathcal{L}^{i,t}_{\ell}=1$ if $S^{\ell}_{i,t}=1$ and $m(\ell)=p$, $\mathcal{L}^{i,t}_{\ell}=2$ if $S^{\ell}_{i,t}=1$ and $m(\ell)=o$. Thus, assume that lobbyists send signals at the beginning of each round $t$ and that the order in which their signals are received by individuals is $Z_{t}\in\text{Perm}\{1,\ldots,L\}$, i.e. an element of the set of all permutations of the indexes of lobbyists (in what follows, we assume that the order according to which lobbyists send their signals is randomly chosen). The weight assigned to the optimistic model deriving from previous interaction needs to be updated for any $z\in\{1,\ldots,L\}$ according to
@@
When lobbyists are active, the timeline of events in each round $t$ becomes: $i)$ each lobbyist sends the signals to the individuals it selected; $ii)$ individual weights are updated according to equation (4); $iii)$ $\iota_{t}$ is independently and uniformly drawn from the population; $iv)$ the set of receivers $J(\iota_{t})$ is created; $v)$ the signal $s_{t}$ is drawn from the distribution ${\bf p}_{\iota_{t},t}$; $vi)$ individual weights are updated according to equation (5).
@@
The second special case we consider is characterised by $N=L=2$, $\tau=1$, $B^{1}=B^{2}=1$, and $\lambda_{i}=\phi_{i}=0\;\forall i\in\{1,2\}$. Without loss of generality, we assume that lobbyist 1 supports the pessimistic model while lobbyist 2 supports the optimistic model. Since both time and budget are unitary (and not sending signals is strictly dominated), each lobbyist has two feasible pure strategies: sending a signal to individual 1 or to individual 2, that is, $S^{\ell}\in\mathcal{S}_{\ell}=\{(0,1),(1,0)\}$. Under the assumptions that the model structure is common knowledge and the lobbyists are rational, we can compute the Nash equilibrium [29] resulting from the strategic interaction of the two lobbyists.
Notice that $\lambda_{i}=\phi_{i}=0\;\forall i\in\{1,2\}$ implies that both individuals update their beliefs according to Bayes rule, hence, the order in which they receive signals does not matter. Thus, to compute final probabilities, we only need to consider the following set of possible received signals: $\Omega=\left\{\emptyset,\{0\},\{1\},\{0,0\},\{0,1\},\{1,1\},\{0,0,1\},\{0,1,1\}\right\}$. That is, since the lobbyists support opposite models and one of the individuals sends one signal to the other, the most extreme cases are the one in which an individual does not receive any signal, and the one in which it receives two (different) signals from the lobbyists and a signal from the other agent. All the other cases in between are possible. Assuming a uniform prior distribution for both individuals, the two agents are ex-ante identical, meaning that the differences in their final probabilities are uniquely generated by the signals they received. Hence, we can simplify the notation calling $w(\omega)$ the weight assigned to the optimistic model by an agent that has received the signals $\omega\in\Omega$. As a direct consequence, we define the probability such an agent assigns to the realization of the event as $p(\omega)=w(\omega)\pi_{o}+(1-w(\omega))\pi_{p}$. Then, calling $t_{s}$, with $s\in\{0,1\}$, the number of times $s$ occurs in $\omega$, one has
@@
In all of our simulations, we consider a complete network of 500 agents. The adjacency matrix $A$ is therefore of size $500\times 500$, with diagonal entries $a_{i,i}=0$, and off-diagonal entries $a_{i,j\neq i}=1$. Each agent is characterised by an individual-specific initial weight $w_{i,0}$ which is randomly drawn from an uniform distribution with support $(0,1)$. The probabilities characterizing the two models (optimistic and pessimistic) are $\pi_{o}=0.01$ and $\pi_{p}=0.99$. We consider three main scenarios. In the “baseline” scenario, the network is left to interact with itself, with no outside interference, i.e. no lobbyist. In the “one-lobbyist” scenario, we introduce a single active lobbyist ($L=1$) which supports the pessimistic model ($m(1)=p$). In the “two-lobbyists scenario”, we introduce two active lobbyists that support competing models ($m(1)=p$, and $m(2)=o$). In each scenario contemplating lobbyists, each one of them is endowed with a budget $B=10,\!000$. For the purpose of our numerical analysis, we introduce the parameter $T\leq\tau$, representing the lobbyists’ time horizon, that is, the time span over which it allocates its budget $B^{\ell}$. Accordingly, we set $S^{\ell}_{i,t}=0\;\forall t\in(T,\tau]$. This constraint allows us to manage the potentially vast space of feasible lobbying strategies and enables analysis in scenarios with large $\tau$. In particular, neglecting the zeros appearing after $T$, we assume that lobbyists dispose of a set of 100 strategy matrices of size $(T\times 500)$ in which the variable indicators for signals are distributed randomly over nodes and iterations, and choose one of those matrices at random at the beginning of each independent simulation run. For most of our experiments, we set $T=100$. This means that, over 100 periods, each lobbyist is able to send 100 signals per period, thus reaching out to 20% of the nodes at each time-step. Moreover, this strategy selection procedure resembles an adaptation of the uniform-probability mixed strategy that emerges as the Nash equilibrium in Example 2.
@@
As shown in equation (3), each $\lambda_{i,t}$ is subject to two different types of bias, captured by the parameters $\lambda_{i}$ and $\phi_{i}$. While $\lambda_{i}$ captures the degree of under-reaction to any new information, $\phi_{i}$ represents the strength of directional motivated reasoning (or confirmation bias), which is signal-dependent. We explore the properties of the model assuming homogeneous biases ($\lambda_{i}=\lambda$ and $\phi_{i}=\phi$ for all individuals $i$) and considering different configurations of the network with respect to these two independent dimensions, i.e. a parameter space generated by $(\lambda,\phi)\in[0,1]\times[0,1]$.

### v3

@@ Abstract
While lobbying has been demonstrated to have an important effect on public opinion and policy making, existing models of opinion formation do not specifically include its effect. In this work we introduce a new model of lobbying-driven opinion influence within opinion dynamics, where lobbyists can implement complex strategies and are characterised by a finite budget. Individuals update their opinions through a learning process resembling Bayes-rule updating but using signals generated by the other agents (a form of social learning), modulated by under-reaction and confirmation bias. We study the model numerically and demonstrate rich dynamics both with and without lobbyists. In the presence of lobbying, we observe two regimes: one in which lobbyists can have full influence on the agent network, and another where the peer-effect generates polarisation. When lobbyists are symmetric, the lobbyist-influence regime is characterised by prolonged opinion oscillations. If lobbyists temporally differentiate their strategies, frontloading is advantageous in the peer-effect regime, whereas backloading is advantageous in the lobbyist-influence regime. These rich dynamics pave the way for studying real lobbying strategies to validate the model in practice.
@@ Introduction
In this paper we focus on the opinion-formation layer of lobbying. In particular, how strategic actors shape public beliefs over time through advocacy, agenda setting and framing, and how grassroots/“outside” mobilization, alongside “inside” access strategies that complement opinion work. These channels are well documented in the political-economy and advocacy literatures and often operate precisely by shifting public attitudes to create political momentum or legitimacy for policy change. Our model abstracts from the policy decision stage and instead endogenizes lobbying-driven opinion influence as costly, time-staged communication under resource constraints. This scope is consistent with empirical evidence that frames and advocacy can materially affect public support and elite incentives [22, 6, 26, 39, 45].
@@
This paper contributes to this growing interdisciplinary literature by introducing a novel model of lobbying influence over opinion dynamics in a boundedly rational social network. In our framework, a population of agents interacts repeatedly to update their subjective beliefs about the likelihood of an uncertain future event, for example, a damaging climate event. Each agent weighs two competing probabilistic models (one “optimistic” and one “pessimistic”) and updates their beliefs through a process resembling Bayes-rule updating but using signals generated by other agents (an extreme form of social learning [15, 4, 2, 25]), modulated by under-reaction [18, 31, 9] and confirmation bias (a form of directional motivated reasoning) [16]. Agents are embedded in a directed network and interact through sequential signal exchanges that propagate beliefs over time. We then endogenize lobbying by introducing external agents (lobbyists) who strategically send costly signals to influence the population’s belief distribution by the end of a finite time horizon. Lobbyists differ in the model they support and are subject to a resource constraint (budget). Their objective is to minimize the distance between the final average belief and their preferred model. Lobbyists select randomized strategies from a constrained set of signalling paths, and their interventions shape both the speed and direction of belief evolution in the network.
@@
Our numerical results reveal rich dynamics. In the absence of lobbying, the network tends to converge toward consensus or, under high confirmation bias, become polarized. A single lobbyist with sufficient resources can dominate belief formation and steer the network toward its preferred model, especially when agents are relatively open to new information. However, when two lobbyists with opposing goals act simultaneously, the system often fails to converge within the active time horizon, resulting in persistent instability and oscillation, especially in regions of the parameter space where both lobbyists are highly effective. We also examine the role of timing in lobbying strategies, showing that frontloading or backloading (i.e. concentrating signals at the beginning or at the end of the interaction time span, respectively) can be advantageous depending on the population’s biases. In the region of the parameter space where lobbying influence is stronger, backloading dominates, whereas in the region where peer effects dominate, frontloading prevails. Overall, our results indicate a complex interplay between cognitive constraints and strategic influence.
Our work relates to the literature on opinion dynamics with stubborn or exogenous influencers, including Acemoglu et al. (2013)[1], that characterizes convergence and fluctuation patterns under linear averaging with stubborn nodes. Our framework differs along three dimensions. First, responsiveness is signal-dependent and state-dependent, capturing under-reaction and directional motivated reasoning (confirmation bias), rather than constant linear weights. Second, influential signals are endogenously generated by budget-constrained lobbyists, arriving as costly, time-staged messages that interact with peer signals. Third, the interaction of these features yields a regime structure with prolonged oscillations under symmetric lobbying which is distinct from fluctuation mechanisms in linear stubborn-agent models.
@@
The remainder of the paper is structured as follows. The next section introduces the baseline opinion dynamics model, describing the agent interaction protocol and belief updating mechanism, and then incorporates lobbying agents and formalizes their strategic behavior, cost constraints, and impact on individual beliefs. The Results section presents the simulation results across three scenarios: no lobbying, single-lobbyist, and dual-lobbyist configurations. It explores the model’s behavior across key parameters (under-reaction and confirmation bias), and performs sensitivity analysis on budget, network size, time horizon and lobbyist timing strategies. We then conclude by summarizing the findings and suggesting avenues for future research.
@@
We consider a social network of individuals who repeatedly interact communicating on the occurrence of an uncertain harmful event. For instance, they can do that by sharing posts, sending messages, and creating contents. We refer to any such item generically as a signal sent by an individual to their neighbors. For simplicity, signals are binary: they either assert that the event will occur or that it will not. Guided by evidence of selective behavior in social networks[49], individuals choose which type of signal to send based on their opinions. For example, an individual who assigns a high probability to the occurrence of the event will be more likely to share a signal affirming that it will occur. Consistent with evidence on social influence and contagion[33, 8, 10, 3], received signals affect the recipients’ opinions. Note that whether signals contain a kernel of truth is not essential for generating opinion change that, in turn, increases the likelihood of sharing similar signals. This is consistent with findings that false news diffuses more than true news on social media[47]. We model opinion formation and updating as a modification of Bayes’ rule: individuals behave as if signals from others are informative about the true data-generating process. At the same time, we allow the updating mechanism to deviate from exact Bayes to capture common processing biases, such as under-reaction[17, 18, 31, 9] and confirmation bias[35, 16]. This setup is also analogous to selective perception[30], a well-documented feature of social-media environments[21, 49]. Hence, our model represents an extreme form of social learning[15, 4, 2, 25], in which individuals do not receive private signals from the true data-generating process but try to learn from what other agents share.
In formal terms, $N$ individuals interact for $\tau$ rounds about the occurrence of an uncertain damaging event after the interaction rounds. Individuals do not know the true probabilities that determine the occurrence of the event, but can rely on two probabilistic models. Those are the probability distributions $(\pi_{o},1-\pi_{o})$ and $(\pi_{p},1-\pi_{p})$, where $\pi_{k}$, with $k\in\{o,p\}$, indicates the probability that the damaging event occurs. The model $(\pi_{o},1-\pi_{o})$ is optimistic while the model $(\pi_{p},1-\pi_{p})$ is pessimistic: $0<\pi_{o}<\pi_{p}<1$ with $\pi_{o}$ close to zero and $\pi_{p}$ close to one.
@@
where $\lambda_{i,t}\in[0,1]$ modulates the effects of cognitive biases of agent $i$ at round $t$ and $w_{i,0}\in(0,1)$ is the initial prior. As pointed out before, our agents build their beliefs in a way that resembles Bayesian updating but is applied in an extreme social-learning setting. They behave as if the observed social signals were i.i.d. draws from the true process, although these signals are noisy observations of other agents’ contemporaneous beliefs. In the absence of new evidence, and under common knowledge of identical $0.5$ priors, Bayesians would keep posteriors at $0.5$. By contrast, our agents treat social signals as informative about the state, departing from proper Bayesian-like learning. We nonetheless interpret $\lambda_{i,t}$ as capturing specific cognitive biases on top of this process. In particular, although proper Bayesian learning cannot be recovered in this social-learning environment, setting $\lambda_{i,t}=0$ makes agent $i$ update by Bayes’ rule and this acts as a benchmark within the proposed setting. If $\lambda_{i,t}=1$ then $i$ keeps the initial combination fixed no matter the sequence of signals it receives. Intermediate values of $\lambda_{i,t}$ indicate that, as $i$ receives a signal, it updates the weights in the right direction but in lower magnitude than what Bayes’ rule would prescribe. Thus, individuals can underreact to signals [18, 31, 9], producing the well-documented pattern in belief updating often referred to as conservatism[17]. We enrich such a behavioural interpretation imposing that $\lambda_{i,t}$ depends upon the type of signal received and the prior. In this way, it captures a form of directional motivated reasoning that, in turn, gives rise to confirmation bias (the tendency of individuals to discard information that contradicts their priors), another common feature in the formation of beliefs [35, 16]. In particular, we assume
@@
with $\phi_{i},\lambda_{i}\in[0,1]$. Note that, our updating rule differs from linear averaging rules commonly studied in stubborn-agent models[1], because the effective step size is endogenous to the signal’s (in)congruence with priors (equation (3)). This feature is central to the regime behavior we document later.
The overall degree of under-reaction is the convex combination of two components. On the one hand, the parameter $\lambda_{i}$ captures the baseline level of under-reaction of agent $i$. On the other hand, the function $|1-s_{t}-w_{i,t-1}|$ introduces directional motivated reasoning or, equivalently for our analysis and interpretation, confirmation bias. The parameter $\phi_{i}$ regulates the strength of the directional motivated reasoning. Suppose, for instance, that directional motivated reasoning is strong ($\phi_{i}\simeq 1$), if an individual has a strong prior in the optimistic model ($w_{i,t-1}\simeq 1$) and receives a signal that favours it with respect to the pessimistic model ($s_{t}=0$) then its overall degree of under-reaction is low: information has a sensible impact on its beliefs. If, instead, the information contradicts its priors ($s_{t}=1$), then the degree of under-reaction is high: information has a negligible effect on its beliefs. A symmetric argument holds if the agent’s priors favour the pessimistic model. If, instead, directional motivated reasoning is weak ($\phi_{i}\simeq 0$), then the individual under-reacts with respect to information because of the effect of $\lambda_{i}$, but the nature of the signal and its relation to the prior have a negligible effect on probability updating. As mentioned above, this setup is also akin to a form of selective perception.
Selective behavior in the generation of signals is modelled assuming that the signal at round $t$ is drawn from the distribution ${\bf p}_{\iota_{t},t}$ (i.e., $\text{Prob}\{s_{t}=1\}=p_{\iota_{t},t}$). Thus, the timeline of events in each round $t$ is as follows: $i)$ $\iota_{t}$ is independently and uniformly drawn from the population; ii)ii) the set of receivers $J(\iota_{t})$ is created; $iii)$ the signal $s_{t}$ is drawn from the distribution ${\bf p}_{\iota_{t},t}$; $iv)$ individual weights are updated according to equation (2).
@@
Lobbyists can be considered external agents with respect to the social network that, in each turn, can pay a cost to send signals to individuals. Their objective is to influence the time $\tau$ distribution of beliefs so that it becomes concentrated on one of the two models. We emphasize that this module captures the opinion-formation channel of lobbying; the policy-adoption stage is not modelled and is treated as exogenous to the opinion dynamics studied here.
@@
Assume that there are $L$ active lobbyists and define $S_{i,t}^{\ell}\in\{0,1\}$ as a variable indicating whether lobbyist $\ell\in\{1,2,\ldots,L\}$ sends a signal to individual $i$ at the beginning of round $t$ ($S^{\ell}_{i,t}=1$) or not ($S^{\ell}_{i,t}=0$). It follows that a pure strategy for lobbyist $\ell$ can be indicated as a matrix ${S}^{\ell}\in\mathcal{S}=\{0,1\}^{N\times\tau}$. Each lobbyist supports one model between the optimist and the pessimist one and the signals it sends are in favour of the model it supports. For instance, a lobbyist supporting the optimistic model will signal 0, that is, the event is not occurring. Define the function that assigns to each lobbyist the model it supports as $m:\{1,2,\ldots,L\}\to\{o,p\},\ell\mapsto m(\ell)$. We assume that signals are costly. In particular, the cost of a signal is fixed to a given level $c>0$ and each lobbyist $\ell$ has a budget $B^{\ell}$. Without loss of generality, we shall set $c=1$ in what follows. That is, $B^{\ell}$ shall be understood as the maximum number of signals lobbyist $\ell$ can send. Notice that this assumption constrains the set of strategies that each lobbyist can actually play, that is, a feasible strategy is such that $\sum_{t=1}^{\tau}\sum_{i=1}^{N}\,S^{\ell}_{i,t}\leq B^{\ell}$. Hence, we define the set of feasible strategies for lobbyist $\ell$ as $\mathcal{S}_{\ell}=\{S^{\ell}\in\mathcal{S}\,|\,\sum_{t=1}^{\tau}\sum_{i=1}^{N}\,S^{\ell}_{i,t}\leq B^{\ell}\}$. We allow lobbyists to randomize their choices. Call $\Delta(\mathcal{S}_{\ell})$ the mixed extension of $\mathcal{S}_{\ell}$ and $\sigma^{\ell}({S})$ the probability that lobbyist $\ell$ attaches to a feasible strategy S{S}, such that $\sigma^{\ell}\in\Delta(\mathcal{S}_{\ell})$ is a feasible mixed strategy for lobbyist $\ell$. Note that, this strategic, budget-constrained sending of costly, time-staged signals differs from exogenous stubborn-node formulations. Indeed, each individual reacts to the signal received in the same way as they do for peer signals. Therefore, now, individual probabilities also evolve as a consequence of the strategies of the lobbyists. In particular, call $S^{-\ell}=(S^{1},\ldots,S^{\ell-1},S^{\ell+1},\ldots,S^{L})\in\mathcal{S}_{1}\times\ldots\times\mathcal{S}_{\ell-1}\times\mathcal{S}_{\ell+1}\times\ldots\times\mathcal{S}_{L}$ a profile of feasible pure strategies for the lobbyists different from $\ell$, such that $p_{i,\tau}(S^{\ell},S^{-\ell})$ indicates the final probability of individual $i$ as a function of the strategies played by all lobbyists. Define the vector of lobbyists’ signals received by agent $i$ at time $t$ as $\mathcal{L}^{i,t}\in\{0,1,2\}^{L}$, where $\mathcal{L}^{i,t}_{\ell}=0$ if $S^{\ell}_{i,t}=0$, $\mathcal{L}^{i,t}_{\ell}=1$ if $S^{\ell}_{i,t}=1$ and $m(\ell)=p$, $\mathcal{L}^{i,t}_{\ell}=2$ if $S^{\ell}_{i,t}=1$ and $m(\ell)=o$. Thus, assume that lobbyists send signals at the beginning of each round $t$ and that the order in which their signals are received by individuals is $Z_{t}\in\text{Perm}\{1,\ldots,L\}$, i.e. an element of the set of all permutations of the indexes of lobbyists (in what follows, we assume that the order according to which lobbyists send their signals is randomly chosen). The weight assigned to the optimistic model deriving from previous interaction needs to be updated for any $z\in\{1,\ldots,L\}$ according to
@@
When lobbyists are active, the timeline of events in each round $t$ becomes: $i)$ each lobbyist sends the signals to the individuals it selected; ii)ii) individual weights are updated according to equation (4); $iii)$ $\iota_{t}$ is independently and uniformly drawn from the population; $iv)$ the set of receivers $J(\iota_{t})$ is created; $v)$ the signal $s_{t}$ is drawn from the distribution ${\bf p}_{\iota_{t},t}$; $vi)$ individual weights are updated according to equation (5).
@@
The second special case we consider is characterized by $N=L=2$, $\tau=1$, $B^{1}=B^{2}=1$, and $\lambda_{i}=\phi_{i}=0\;\forall i\in\{1,2\}$. Without loss of generality, we assume that lobbyist 1 supports the pessimistic model while lobbyist 2 supports the optimistic model. Since both time and budget are unitary (and not sending signals is strictly dominated), each lobbyist has two feasible pure strategies: sending a signal to individual 1 or to individual 2, that is, $S^{\ell}\in\mathcal{S}_{\ell}=\{(0,1),(1,0)\}$. Under the assumptions that the model structure is common knowledge and the lobbyists are rational, we can compute the Nash equilibrium [34] resulting from the strategic interaction of the two lobbyists.
Notice that with $\lambda_{i}=\phi_{i}=0\;\forall i\in\{1,2\}$ individuals follow Bayes-rule updating under the extreme social learning framework that characterizes our model. In this one-period, signal-symmetric case, the order of receipt is immaterial for the final posteriors. Thus, to compute final probabilities, we only need to consider the following set of possible received signals:
$\Omega=\left\{\emptyset,\{0\},\{1\},\{0,0\},\{0,1\},\{1,1\},\{0,0,1\},\{0,1,1\}\right\}.$
That is, since the lobbyists support opposite models and one of the individuals sends one signal to the other, the most extreme cases are the one in which an individual does not receive any signal, and the one in which it receives two (different) signals from the lobbyists and a signal from the other agent. All the other cases in between are possible. Assuming a uniform prior distribution for both individuals, the two agents are ex-ante identical, meaning that the differences in their final probabilities are uniquely generated by the signals they received. Hence, we can simplify the notation calling $w(\omega)$ the weight assigned to the optimistic model by an agent that has received the signals $\omega\in\Omega$. As a direct consequence, we define the probability such an agent assigns to the realization of the event as $p(\omega)=w(\omega)\pi_{o}+(1-w(\omega))\pi_{p}$. Then, calling $t_{s}$, with $s\in\{0,1\}$, the number of times $s$ occurs in $\omega$, one has
@@
In all of our simulations, we consider a complete network of 500 agents. The adjacency matrix $A$ is therefore of size $500\times 500$, with diagonal entries $a_{i,i}=0$, and off-diagonal entries $a_{i,j\neq i}=1$. Each agent is characterised by an individual-specific initial weight $w_{i,0}$ which is randomly drawn from an uniform distribution with support $(0,1)$. The probabilities characterizing the two models (optimistic and pessimistic) are $\pi_{o}=0.01$ and $\pi_{p}=0.99$. We consider three main scenarios. In the “baseline” scenario, the network is left to interact with itself, with no outside interference, i.e. no lobbyist. In the “one-lobbyist” scenario, we introduce a single active lobbyist ($L=1$) which supports the pessimistic model ($m(1)=p$). In the “two-lobbyists scenario”, we introduce two active lobbyists that support competing models ($m(1)=p$, and $m(2)=o$). In each scenario contemplating lobbyists, each one of them is endowed with a budget B=10, [^0] B=10,\!000. For the purpose of our numerical analysis, we introduce the parameter $T\leq\tau$, representing the lobbyists’ time horizon, that is, the time span over which it allocates its budget $B^{\ell}$. Accordingly, we set $S^{\ell}_{i,t}=0\;\forall t\in(T,\tau]$. This constraint allows us to manage the potentially vast space of feasible lobbying strategies and enables analysis in scenarios with large $\tau$. In particular, neglecting the zeros appearing after $T$, we assume that lobbyists dispose of a set of 100 strategy matrices of size $(T\times 500)$ in which the variable indicators for signals are distributed randomly over nodes and iterations, and choose one of those matrices at random at the beginning of each independent simulation run. For most of our experiments, we set $T=100$. This means that, over 100 periods, each lobbyist is able to send 100 signals per period, thus reaching out to 20% of the nodes at each time-step. Moreover, this strategy selection procedure resembles an adaptation of the uniform-probability mixed strategy that emerges as the Nash equilibrium in Example 2. Throughout, we interpret results as effects on lobbying-driven opinion influence; any policy-adoption mapping is intentionally left outside the model.
Parameter Description Baseline value
NN Number of individuals in the network 500
$T$ Number of interaction rounds 100
$\pi_{o}$ Probability of damaging event under optimistic model 0.01
$\pi_{p}$ Probability of damaging event under pessimistic model 0.99
$\phi$ Strength of directional motivated reasoning/confirmation bias $[0,1]$
$\lambda$ Degree of baseline under-reaction $[0,1]$
$B$ Budget of each lobbyist 10000
$c$ Cost per signal sent by each lobbyist 1
Table 1: Summary of model’s parameters, their descriptions, and their baseline numerical values.
@@
As shown in equation (3), each $\lambda_{i,t}$ is subject to two different types of bias, captured by the parameters $\lambda_{i}$ and $\phi_{i}$. While $\lambda_{i}$ captures the degree of under-reaction to any new information, $\phi_{i}$ represents the strength of directional motivated reasoning (or confirmation bias), which is signal-dependent. We explore the properties of the model assuming homogeneous biases ($\lambda_{i}=\lambda$ and $\phi_{i}=\phi$ for all individuals $i$) and considering different configurations of the network with respect to these two independent dimensions, i.e. a parameter space generated by $(\lambda,\phi)\in[0,1]\times[0,1]$. For reference, Table 1 summarizes the model parameters, their descriptions, and the baseline values.